OpenStudy (anonymous):

what is the derivative of cosh(1/(x+e^x))?

OpenStudy (anonymous):

not to bad but kind of ugly chain rule for this one

OpenStudy (anonymous):

what is the derivative of \(\cosh(x)\) ?

OpenStudy (anonymous):

-sin h (x)

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

this is what i got... -sinh(1/(x+e^x) (1x+e^x)^-2 + (1+e^x)

OpenStudy (anonymous):

you are getting that confused with the derivative of sine being minus cosine

OpenStudy (anonymous):

i thought the derivative of sin is just cos

OpenStudy (anonymous):

the derivative of \(\cosh(x)\) is \(\sinh(x)\)

OpenStudy (anonymous):

and the derivative of cos is negative sin

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yeah i wrote that backwards what i meant is you are getting it confused with the derivative of cosine being minus sine

OpenStudy (anonymous):

woahhh i didn't know that

OpenStudy (anonymous):

so is it just what i wrote negated.

OpenStudy (anonymous):

sinh(1/(x+e^x))(x+e^x)^-2+1+e^x?

OpenStudy (anonymous):

then you need the chain rule

OpenStudy (anonymous):

you get \[\sinh(\frac{1}{x+e^x})\times \frac{d}{dx}[\frac{1}{x+e^x}]\]